OpenStudy (anonymous):

Solve the following equation: 5(2x - 7) = 4x + 13

OpenStudy (anonymous):

start by multiplying out the left side. It is distributive, so multiply the 5 by both things in the parentheses

OpenStudy (anonymous):

8

OpenStudy (anonymous):

great name!

OpenStudy (anonymous):

10x-35=4x+13 10x-4x-35=13 6x-35=13 6x=13+35 6x=48 x=48/6 x=8 ok

OpenStudy (anonymous):

ok man?